Circular No. 8615

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005gk AND 2005gl Further to IAUC 8610, T. Puckett reports the discovery of two appar-

ent supernovae on unfiltered CCD frames taken with the 0.60-m automated supernova patrol telescope — 2005gk in collaboration with D. George and 2005gl with P. Ceravolo.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005gk Oct. 4.31 3^{h}08^{m}16.^{s}05 − 4^{o}12′00′′.2 18.5 5′′.5 E, 3′′.5 N
2005gl Oct. 5.18 0 49 50.02 +32 16 56.8 18.2 29′′.8 E, 16′′.7 N

S. Nakano, Sumoto, Japan, reports the indepedent discovery of 2005gl by Y. Sano (Nayoro, Hokkaido) on thirty unfiltered CCD frames taken around Oct. 12.57 UT with a 0.28-m f/6.1 Schmidt-Cassegrain reflector, providing position end figures 50^{s}.11, 55′′.9. Nakano adds that Y. Kushida (Yatsugatake South Base Observatory, 0.40-m f/10.0 Schmidt-Cassegrain reflector + unfiltered CCD) provides position end figures 50^{s}.12, 56′′.1 for 2005gl from an exposure taken by R. Kushida. Additional approximate CCD magnitudes: SN 2005gk, 2004 Oct. 13 and 27, [20.0 (Puckett); 2005 Oct. 9.19, 18.5 (Puckett). SN 2005gl in NGC 266, Sept. 3 and 10, [20.0 (Puckett); Oct. 5, [18 (Sano); 12.14, 16.8 (D. George, Ottawa, ON, 0.40-m reflector; via Puckett); 12.57, 16.7 (Sano); 12.632, 17.0 (Kushida).

M. Ganeshalingam, M. R. Moore, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330– 1000 nm), obtained on Oct. 11 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2005gk is of type II. The P-Cyg profile of the broad Hα line is dominated by the emission component.

COMETS C/2005 P4, 2005 P5, 2005 P6, 2005 Q5 (SOHO) The following four tailless Kreutz-sungrazing comets, found on SOHO

website images (cf. IAUC 8604), were all stellar in appearance. C/2005 P4 reached mag 7.6 at 12.2R_{⊙} on Aug. 11.611 UT. C/2005 P5 reached mag ∼ 8 at ∼ 11R_{⊙}. C/2005 P6 reached mag 7.7 at 11.9R_{⊙} on Aug. 14.321. C/2005 Q5 reached mag 7.3 at 10.8R_{⊙} on Aug. 18.321.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 P4 Aug. 11.071 9^{h}01.^{m}4 +13^{o}35′ C3 TH 2005-T55
C/2005 P5  12.638 9 09.7 +13 49 C3 HS 2005-T55
C/2005 P6  14.029 9 16.2 +13 08 C3 HS 2005-T55
C/2005 Q5  17.821 9 27.6 +12 00 C3 SF 2005-T55
2005 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green